Consider a normally distributed population with a mean of 4 and a variance of 16. If a random sample of size 16 is drawn, what is the probability that the sample mean is at most 4?


1
Expert's answer
2021-09-07T19:02:23-0400

Solution:

Given, "\mu=4,\sigma^2=16,n=16"

"X\sim(\mu,\sigma)"

Then, "\bar X\sim(\mu,\sigma/\sqrt n)"

"P(\bar X\le4)=P(z\le \dfrac{4-4}{4/4})
\\=P(z\le 0)
\\=0.5"